# Self-Supervised Interpretable Sensorimotor Learning via Latent Functional Modularity

### v1

@@ Introduction
Considering that the latent decision contains task-oriented features, we decode the decision vector $h^{d}$ into an understandable, task-wise probability score vector to facilitate the behavioral interpretation of our network. We employ a multiclass linear Support Vector Machine (SVM) classifier (Suthaharan 2016) that is computationally efficient and less prone to overfitting. Utilizing sample decisions and their corresponding task labels $(h^{d}_{i},y_{i})$, linear SVM is designed to learn binary classification through the following optimization  (Tang 2013):
@@
where $w$ is the weight vector, $b$ is the bias term, and $C$ is the regularization parameter. The linear SVM is extended to multiclass classification using the one-vs-rest scheme (Tang 2013). This adaptation enables SVMs to maximize the margin between input data belonging to different classes. After training the multiclass linear SVMs, we transform their output into a posterior probability score vector $P(y_{i}\!=k|h^{d}_{i})$ for each class $k$, using a calibration method (Niculescu-Mizil and Caruana 2005):
@@
We carry out behavior interpretation in a post-hoc manner. Initially, we generate sample latent decisions for each task that human engineers aim to interpret, using the trained modular network. Subsequently, we train the multiclass SVMs, along with parameters $E_{k}$ and $F_{k}$ for the calibration method. Finally, during sensorimotor inferencing, we transform MoNet’s latent decisions into probability score vectors using Eq. 12. This approach allows us to interpret the end-to-end model without sacrificing sensorimotor performance.
Our approach is comparable to concept-based interpretation methods in explainable artificial intelligence (Ghorbani et al. 2019). These studies focus on understanding how high-level concepts are represented and utilized by models in decision-making. In our case, the concept vector corresponds to the latent decision that encapsulates the driving situation. Consequently, to interpret the decision intent during the sensorimotor process, we quantify the alignment of a given decision vector with the specific tasks’ concept (driving situation). This is performed by decoding the latent decision into the understandable probability score vector.
Experiments
@@
Fig. 3 shows the overall setup for hardware, environment, and scenarios in this work. We apply MoNet on a wheeled, car-like platform, modeled after the F1TENTH vehicle (O’Kelly et al. 2020). Our platform consists of a 1/10 scale racing car chassis (TT-02) equipped with an embedded computer (Jetson Xavier NX) and a controller (Arduino). The Xavier NX receives front camera images and range measurements from sensors mounted on the platform. The range measurements are utilized to estimate the current pose of the ego vehicle and compute ego-centric coarse topology map (Amini et al. 2019). Detailed hardware setup and the coarse map processing are provided in the appendix.
@@
Our platform performs visual autonomous navigation with multiple driving tasks such as straight (ST), left-turn (LT), right-turn (RT), and collision avoidance (CA). Data collection is carried out under controlled conditions in two indoor environments: Corridor Environment 1 (Env. 1, $71m\times 16m$) and Environment 2 (Env. 2, $88m\times 35m$). Box-shaped obstacles are randomly positioned within specific areas in these environments. The training dataset comprises data from scenarios that feature either a single obstacle or no obstacles. However, scenarios involving multiple obstacles are introduced as new, unseen challenges during the evaluation phases. Our method is evaluated in Env. 1, characterized by more frequent intersection situations during autonomous navigation. For further details on data collection and processing (see the appendix for details).
@@
Quantitative Evaluation
Sensorimotor Performance
We evaluate the sensorimotor performance by measuring the success rate for each driving task in the evaluation environment. In addition to our method, we have implemented ViTNet, a baseline model designed as a conventional Vision Transformer-based end-to-end architecture without the planning module ($\mathcal{Q}$) from the MoNet’s architecture. Under the same hardware and environmental conditions, each model performed 16 episodes in the real-world environment, totaling 143 driving tasks. Table 1 summarizes the performance results. These results demonstrate that our model exhibits stronger generalization ability across multiple sensorimotor tasks compared to the baseline model. Both models show safe navigation performance in straight driving scenarios. However, ViTNet often struggles to overcome unseen obstacle scenarios and particularly fails in turning right at intersections, where it records its lowest success rate of 63%. Although there was a situation where our model had a mild touch with a wall while avoiding cluttered obstacles, MoNet succeeded in all trials of navigating intersections and generally performed well in obstacle avoidance scenarios.
Planning Performance
To assess the planning-level performance of the end-to-end network, we quantify the task specificity during sensorimotor inference using a t-SNE map (Van der Maaten and Hinton 2008) and a Representational Similarity Matrix (RSM) (Popal, Wang, and Olson 2019). In consideration of the data distribution, we sampled 318, 64, 59, 59, and 67 pieces of data, respectively, for the five tasks (ST, SI, LT, RT, CA) for these assessments.
The t-SNE visualization (Fig. 4 (A)) demonstrates that our planning module generates distinct and well-structured decisions in the latent space for various tasks. It effectively differentiates between the LT and RT tasks from the ST scenario and recognizes the directional variations in navigating intersections. The decisions for CA are positioned between the ST, LT, and RT clusters, indicating a need for moderate planning that involves both intersection-turning and corridor-following behaviors in collision avoidance scenarios. Furthermore, the data for SI exhibits a high similarity to that of ST, reflecting similar driving contexts, despite their differing inputs from topological maps.
For a more quantitative analysis of similarity in latent planning, we further examine the RSM of the learned decisions (Fig. 4 (B)). We measure the cosine similarity between the latent decisions of each task using the average linkage method (Nielsen and Nielsen 2016). The similarities among the five task classes are then normalized to a range of [0, 1] through a row-wise softmax operation. The results demonstrate that our method effectively differentiates between various tasks while clustering similar driving situations. The matrix shows strong diagonal values, indicating that the latent decisions effectively distinguish various driving tasks based on contextual features derived from sensory data, without task-level inputs. Specifically, the decisions related to navigating intersections (LT, RT) and avoiding collisions (CA) are clearly differentiated from other tasks, ensuring minimal confusion. Additionally, the decisions from similar contexts, such as ST and SI, although showing higher similarities than the average value of 0.2, remain distinguishable from each other. This analysis provides the user with a clear understanding of the network’s performance in discriminating between different tasks. It also allows for the consolidation of similar task classes (ST and SI in our case), which increases the efficiency of decision interpretation.
Figure 5: Quantitative results showing given sensory inputs, latent decisions, decoded interpretable decisions, entropy, and control output during an autonomous navigation episode.
Learning Curves
We evaluate the learning curves of various baseline models based on their control and planning performance (Fig. 4 (C)). For sensorimotor control performance, we calculate the L1 loss using the validation dataset. To assess planning-level performance, we compute a similarity score, which is the sum of the diagonal values in the RSM results. In addition to our method, we configure ViTNet, which comprises only perception and control modules, to investigate the necessity of the planning module. For ViTNet’s similarity score computation, we select the perceptual feature ($z^{p}$) and the control-level hidden states ($z^{c}$), where $z^{c}$ is the output of Eq. 5, computed without the top-down decision process. Additionally, we introduce another version of MoNet (MoNet w/o $L_{\text{CGC}}$), trained without the contrastive loss $L_{\text{CGC}}$, to analyze the effectiveness of the self-supervised scheme. Our approach achieves notable improvement in top-down latent planning over other models, without compromising sensorimotor learning capabilities. The L1 loss curves show minimal changes with the addition of an extra planning module or a contrastive learning scheme. Meanwhile, in the similarity score curves, our method outperforms other approaches, demonstrating 11%-47% higher final performance. MoNet (w/o $L_{\text{CGC}}$), which only includes a planning module, does not exhibit significant discrimination performance (1.00) in the latent space. In contrast, the latent decision ($h^{d}$) of our modular network (MoNet), utilizing the self-supervised contrastive scheme, achieves the highest terminal score (1.47), outperforming the perception-level (1.32) and control-level (1.13) hidden features. This indicates that our method substantially enhances the task-specificity of end-to-end inferences without affecting policy learning. Moreover, it reveals that our latent decision more effectively embeds contextual characteristics compared to perceptual or low-level control features, thereby improving its suitability for behavioral interpretation.
Analysis of Interpretability
@@
We investigate the interpretability and transparency of our model while performing end-to-end sensorimotor processing by decoding the top-down latent decisions. Fig. 5 illustrates the quantitative results, including latent decisions, decoded interpretable decisions, and control output, during an autonomous navigation episode encompassing multiple tasks. Since the decoded decisions are represented as probabilistic score vectors, we further compute the entropy of these decisions. This entropy represents the confidence level of the internal decision-making during sensorimotor processing. The results show that our method can provide interpretable sensorimotor processes through decoded decisions that validly reflect the driving situation based on sensory inputs. In the early phase, our robot followed a straight corridor using minimal steering control, demonstrating strong probability scores for the task decision ST. However, when navigating intersections, our model produced latent decisions that were decoded as high scores for corresponding turns (LT, RT), necessitating large steering angle commands. In the case of approaching a wall or obstacle, our network generates different patterns (CA) of neural activations, resulting in unique decision responses compared to straight driving (ST) and turns (LT, RT).
@@
Moreover, by analyzing the entropy of the decoded probability score vector, we can assess the confidence level of internal decision-making during end-to-end control. Whenever the robot needed to alter its current driving decisions, such as when approaching intersections or obstacles, the entropy of the decision increased to more than 1.0, indicating mid-level entropy values. Since latent decision-making is the causal process leading to robot control, our method can provide the internal confidence of the end-to-end inference prior to executing the robot’s control actions. This shows the significant interpretability and transparency of our model from a behavioral perspective in practical applications.
@@
We further delve into the perceptual and behavioral interpretation across various driving tasks and include these results in the appendix for brevity.
Conclusion
@@
We introduced MoNet, a modular end-to-end network for self-supervised and interpretable sensorimotor learning. The network leverages functional modularity to enable a novel cognition-guided contrastive scheme. This scheme allows the network to learn task-specific sensorimotor control without task-level supervision. Furthermore, our network incorporates a self-attention mechanism and an internal decision process, both of which can be decoded into a spatial saliency map and an explainable decision. In real-world autonomous navigation, our model demonstrates effective sensorimotor performance with interpretability among multiple tasks.
@@
Our approach to interpretable sensorimotor learning with functional modularity offers several advantages for the use of end-to-end networks. Firstly, it enables more reliable and less uncertain end-to-end processes in robotics. Our method allows human engineers to comprehend the network’s intent and the rationale behind specific control outputs from perspectives beyond control-level observation, including perception and planning. Such enhancement is particularly valuable in real-world deployments where safety is critical. Secondly, our approach facilitates the integration of learning-based, black-box modules with nonlearning-based, white-box ones into a hybrid architecture. By leveraging decoded interpretable decisions from our modular network, it becomes feasible to conditionally apply either network-based policies or conventional controllers during deployment. We hope that our work contributes to integrating robotic sensorimotor processes with explainable artificial intelligence. We plan to validate our method across a more diverse range of robotic platforms and scenarios.
Acknowledgments
This work was supported by the Technology Innovation Program (RS-2023-00256794, Development of drone-robot cooperative multimodal delivery technology for cargo with a maximum weight of 40kg in urban areas) funded By the Ministry of Trade, Industry & Energy(MOTIE, Korea).
Supplementary Materials
@@ Acknowledgments
A demo video, codes, and dataset for quantitative results and interpretation are available at https://drive.google.com/drive/folders/1zREu8PdCf1D6hre06dnSpni8FPGwU6vT.
Perception Module in Details
@@
The perception module utilizes the Transformer encoder to generate a saliency map, which is then integrated with the global context of the input image $I$ through the self-attention mechanism. Following the description in (Dosovitskiy et al. 2020), the encoder network comprises: 1) a multi-head self-attention block (MSA), and 2) an MLP block, both equipped with layer normalizations (LN) and residual connections. After feature extraction by the CNN block (Eq. 13), the input embedding undergoes preprocessing (Eq. 14-15) before being fed into the Transformer encoder process (Eq. 16-18).
@@
Hardware System and Coarse Topology Map
Our platform consists of a 1/10 scale racing car chassis equipped with an embedded computer, Jetson Xavier NX, and a controller, Arduino. The Xavier NX receives front camera images with the Realsense D435i camera sensor and acquires range measurements using a 2D LiDAR sensor (Hokuyo UST-20LX). These measurements are utilized to estimate the current pose of the ego vehicle through onboard localization (Hess et al. 2016) in GPS-denied indoor environments. The Xavier NX then computes an ego-centric coarse topology map (Amini et al. 2019), which includes a highlighted routed map alongside an unrouted map, based on the ego vehicle’s pose and a globally routed path. This path is planned using the Dijkstra algorithm, utilizing a sparse topological roadmap of the indoor corridor environments. The Arduino receives commands from either the Xavier NX or a human driver, converting them into PWM signals for the steering and speed control motors of the platform.
Data Collection and Processing
While collecting data, we record camera images, topology maps, and corresponding command signals for steering and throttle control from the human driver. These control signals are normalized to a range of [-1, +1]. We collect data for a total of 2 hours, amounting to 88,326 pairs of sensory input and labels in the environments of Env. 1 and Env. 2. The data is split into training and validation sets at a ratio of $80:20$. The camera image is cropped to a size of $440\times 240$ pixels and then resized to $224\times 224$ pixels for use in our network. For data augmentation, we apply random image shifts and corresponding steering angle adjustments, as outlined in (Bojarski et al. 2016).
Perceptual and Behavioral Interpretability

### v2

@@ 3.4 Interpretation Details
Considering that the latent decision contains task-oriented features, we decode the decision vector $h^{d}$ into an understandable, task-wise probability score vector to facilitate the behavioral interpretation of our network. We employ a multiclass linear Support Vector Machine (SVM) classifier (Suthaharan, 2016) that is computationally efficient and less prone to overfitting. Utilizing sample decisions and their corresponding task labels $(h^{d}_{i},y_{i})$, linear SVM is designed to learn binary classification through the following optimization (Tang, 2013):
@@
where $w$ is the weight vector, $b$ is the bias, and $C$ is the regularization parameter. The SVM is extended to multiclass classification using the one-vs-rest scheme. This adaptation enables SVMs to maximize the margin between input data belonging to different classes. After training the multiclass SVMs, we transform their output into a posterior probability score vector $P(y_{i}\!=k|h^{d}_{i})$ for each class $k$, using a calibration method (Niculescu-Mizil & Caruana, 2005):
@@
We carry out behavior interpretation in a post-hoc manner. Initially, we generate sample latent decisions for each task that human engineers aim to interpret, using the trained modular network. Subsequently, we train the multiclass SVMs, along with parameters $E_{k}$ and $F_{k}$ for the calibration method. Finally, during sensorimotor inferencing, we transform MoNet’s latent decisions into score vectors using Eq. 12 (Fig. 4). This approach allows us to interpret the end-to-end model without sacrificing sensorimotor performance.
Our approach is comparable to concept-based interpretation methods in explainable artificial intelligence (Ghorbani et al., 2019). These studies focus on understanding how high-level concepts are represented and utilized by models in decision-making. In our case, the concept vector corresponds to the latent decision that encapsulates the driving situation. Consequently, to interpret the decision intent during the sensorimotor process, we quantify the alignment of a given decision vector with the specific tasks’ concept (driving situation). This is performed by decoding the latent decision into the understandable posterior probabilities.
4 Experiments
@@ 4.1 Experimental Setup
Fig. 5 shows the overall setup for hardware, environment, and scenarios in this work. We apply MoNet on a wheeled, car-like platform, modeled after the F1TENTH vehicle (O’Kelly et al., 2020). Our platform consists of a 1/10 scale racing car chassis (TT-02) equipped with an embedded computer (Jetson Xavier NX) and a controller (Arduino). The Xavier NX receives front camera images and range measurements from sensors mounted on the platform. The range measurements are utilized to estimate the current pose of the ego vehicle and compute ego-centric coarse topology map (Amini et al., 2019). Detailed hardware setup and the coarse map processing are provided in Appendix C.2.
@@
Our platform performs visual autonomous navigation with multiple driving tasks such as straight (ST), straight-intersection (SI), left-turn (LT), right-turn (RT), and collision avoidance (CA). Data collection is carried out under controlled conditions in two indoor environments: Corridor Environment 1 (Env. 1, $71m\times 16m$) and Environment 2 (Env. 2, $88m\times 35m$). Box-shaped obstacles are randomly positioned within specific areas in these environments. The training dataset comprises data from scenarios that feature either a single obstacle or no obstacles. However, scenarios involving multiple obstacles are introduced as new, unseen challenges during the evaluation phases. Our method is evaluated in Env. 1, characterized by more frequent intersection situations during autonomous navigation. For further details on data collection and processing, we refer to Appendix C.3.
4.2 Quantitative Evaluation
Baseline Models
In addition to our method, we have implemented ViTNet, a baseline model designed as a Vision Transformer-based end-to-end architecture comprising only perception ($\mathcal{P}$) and control ($\mathcal{R}$) modules. This design allows us to investigate the necessity of the planning module ($\mathcal{Q}$). For comparison with the latent decision, we select the perceptual ($z^{p}$) and the control-level hidden features ($z^{c}$) of ViTNet. Here, $z^{c}$ is the output of Eq. 5, computed without involving the internal decision process. Additionally, we introduce MoNet-based methods, MoNet-MUL, MoNet-Iden, and MoNet-NoLGC, for the ablation study. To analyze latent decision computation, MoNet-MUL is configured to perform element-wise multiplication instead of an additive process in Eq.5. MoNet-Iden, whose planning module acts as an identity function, is developed to assess the impact of neural processing within the planning module. MoNet-NoLGC is trained without the LGC loss function to investigate the impact of our self-supervised contrastive learning approach on the task specificity of the network.
Planning Performance
To assess the planning-level performance of the end-to-end network, we quantify the task specificity during sensorimotor inference using a t-SNE map (Van der Maaten & Hinton, 2008) and a Representational Similarity Matrix (RSM) (Popal et al., 2019). These analyses provide the user with a clear understanding of the network’s performance in discriminating between different tasks. In consideration of the data distribution, we sampled 318, 64, 59, 59, and 67 pieces of data, respectively, for the five tasks (ST, SI, LT, RT, CA) for these assessments.
The t-SNE visualization (Fig. 6 (A)) demonstrates that our planning module generates distinct and well-structured decisions in the latent space for various tasks. It effectively differentiates between the LT and RT tasks from the ST scenario and recognizes the directional variations in navigating intersections. The decisions for CA are positioned between the ST, LT, and RT clusters, indicating a need for moderate planning that involves both intersection-turning and corridor-following behaviors in collision avoidance scenarios. Furthermore, the data for SI exhibits a high similarity to that of ST, reflecting similar driving contexts, despite their differing inputs from topological maps. This result highlights that our network is adept at capturing the common driving context found in straight driving, whether it occurs in corridors or intersections. Given these findings, and to ensure a clearer distinction of task classes, we have decided to classify ST and SI as the same task, designated as ST, in the experiments discussed later in this manuscript.
For a more quantitative analysis of latent planning, we further examine the RSM of the learned decisions across different baseline models (Fig. 6 (B)). We measure the cosine similarity between the latent decisions of each task using the average linkage method. The similarities among the four classes are then normalized to a range of [0, 1] through a row-wise softmax operation. The results demonstrate that MoNet effectively differentiates between various tasks while clustering similar situations. The matrix shows strong diagonal values, indicating that the latent decisions effectively distinguish various driving tasks based on contextual features derived from sensory data, without task-level inputs. While the three models—MoNet-MUL, MoNet-Iden, and ViTNet($z^{p}$)—show high similarity values sufficient to separate multiple tasks, they struggle to distinguish between the directional characteristics of LT and RT. ViTNet($z^{c}$), which relies solely on control-level features, fails to represent discriminative task specificity among the four tasks. Interestingly, in the absence of our LGC loss, the modular network MoNet-NoLGC also fails to learn task-specific features, reminiscent of the ’collapse’ issue mentioned in (Mittal et al., 2022). This underscores the effect of our LGC loss in preventing collapse, significantly increasing the specialization of latent planning within the end-to-end network, even in a self-supervised manner.
Learning Curves
We evaluate the learning curves of the baseline models based on their control and planning performance (Fig. 6 (C)). For control performance, we calculate the L1 loss using the validation dataset. To assess planning-level performance, we compute a similarity score, which is the sum of the diagonal values in the RSM results. We skip the model MoNet-NoLGC because it does not show a meaningful similarity score compared to other baseline models (near 1.0). Our approach achieves notable improvement in latent planning over other models, without compromising sensorimotor learning capabilities. The L1 loss curves show minimal changes with the addition of an extra planning module or a contrastive learning scheme. This indicates that our method substantially enhances the task-specificity of end-to-end inferences without affecting policy learning. Meanwhile, in the similarity score curves, our method outperforms other approaches, demonstrating 7%-28% higher final performance than ViTNet-based methods. The latent decision ($h^{d}$) of MoNet, utilizing the self-supervised contrastive scheme, achieves a terminal score of 1.47, outperforming the perception-level (1.37) and control-level (1.15) hidden features from ViTNet. Even when utilizing the identity function, the latent decision-making of MoNet-Iden shows better improvement in task specificity (1.29) compared to the control-level features. This reveals that our latent decision-based approach embeds contextual characteristics more effectively compared to perceptual or low-level control features. Although MoNet-MUL achieves the highest terminal score (1.58), based on the results of the RSM analysis, we have chosen MoNet with the additive process as our primary approach. This approach is selected for its ability to clearly address the four multiple driving tasks.
Figure 7: Quantitative results showing given sensory inputs (front camera images with topological maps), latent decisions, decoded interpretable decisions, entropy, and control output during an autonomous navigation episode.
Sensorimotor Performance
We evaluate the sensorimotor performance of MoNet by measuring the success rate of each task within the evaluation environment, comparing it with ViTNet, which features a perception-control-based end-to-end architecture. Under the same hardware and environmental conditions, each model performed 16 episodes in the real-world environment, totaling 143 driving tasks. Table 1 summarizes the performance results. These results demonstrate that our model exhibits stronger generalization ability across multiple sensorimotor tasks compared to the baseline model. Both models show safe navigation performance in straight driving scenarios. However, ViTNet often struggles to overcome unseen obstacle scenarios and particularly fails in turning right at intersections, where it records its lowest success rate of 63%. Although there was a situation where our model had a mild touch with a wall while avoiding cluttered obstacles, MoNet succeeded in all trials of navigating intersections and generally performed well in obstacle avoidance scenarios.
@@ 4.2 Quantitative Evaluation
4.3 Analysis of Interpretability
@@
We investigate the interpretability and transparency of our model while performing end-to-end sensorimotor processing by decoding the top-down latent decisions. Fig. 7 illustrates the quantitative results, including latent decisions, decoded interpretable decisions, and control output, during an autonomous navigation episode encompassing multiple tasks. Since the decoded decisions are represented as probabilistic score vectors, we further compute the entropy of these decisions. This entropy represents the confidence level of the internal decision-making during end-to-end processing. The results show that our method can provide interpretable sensorimotor processes through decoded decisions that validly reflect the driving situation based on sensory inputs. In the early phase, our robot followed a straight corridor using minimal steering control, demonstrating strong probability scores for the task decision ST. However, when navigating intersections, our model produced latent decisions that were decoded as high scores for corresponding turns (LT, RT), necessitating large steering commands. In the case of approaching a wall or obstacle, our network generates different patterns of neural activations (CA), resulting in unique decision responses compared to straight driving and turns (LT, RT). These results highlight that our approach enables a novel investigation of latent decision transitions during end-to-end inferences, thereby enhancing the transparency of online sensorimotor processes.
@@
Moreover, by analyzing the entropy of the probability score vector, we can assess the confidence level of internal decision-making during end-to-end control. Whenever the robot needed to alter its current driving decisions, such as when approaching intersections or obstacles, the entropy of the decision increased to more than 1.0, indicating mid-level uncertainty values. Since latent decision-making is the causal process leading to robot control, our method can provide the internal confidence of the end-to-end inference prior to executing the robot’s actions. This shows the significant interpretability of our model from a behavioral perspective in practical, real-world applications.
@@
We further delve into perceptual and behavioral interpretation across various tasks by visualizing spatial saliency maps and interpretable decisions. We include these supplementary results in Appendix C.4 for brevity.
5 Conclusion and Limitations
@@
We introduced MoNet, a modular network for self-supervised and interpretable end-to-end learning. Our method leverages functional modularity to enable a novel latent-guided contrastive learning scheme. This scheme allows the network to learn task-specific sensorimotor control without the need for task-level supervision. Furthermore, our network incorporates a self-attention mechanism and an internal decision process, both of which can be decoded into a spatial saliency map and an explainable decision. In real-world autonomous navigation, our model demonstrates effective sensorimotor performance with interpretability among multiple driving tasks.
@@
Our approach to interpretable end-to-end learning with functional modularity offers several advantages for the use of end-to-end network architectures. Firstly, it enables more reliable and less uncertain end-to-end processes in robotics. Our method allows human engineers to comprehend the network’s intent and the rationale behind control outputs from perspectives beyond control-level observation, including perception and planning. Such enhancement is particularly valuable in real-world deployments where safety is critical. Secondly, our approach facilitates the integration of learning-based, black-box modules with nonlearning-based, white-box ones into a hybrid architecture. By leveraging decoded interpretable decisions from our modular network, it becomes feasible to conditionally apply either network-based policies or conventional controllers during deployment. We hope our work contributes to integrating explainable artificial intelligence with end-to-end learning schemes, thereby enhancing the interpretability and transparency of learning-based robotic applications.
While MoNet shows promising results in real-world indoor environments, our method needs further extension to navigate more complex and dynamic environments, such as outdoor scenarios. Since these scenarios contain dynamic and varying features (e.g., moving objects, brightness), temporal features are crucial. An avenue for future work is to incorporate temporal network layers, such as LSTM, into our Vision Transformer-based perception module to learn temporally consistent features in dynamic driving scenes. We believe such a spatio-temporal module will enable our method to capture distinct task-level features with temporal consistency from a perceptual feature perspective. This will be one of the primary focuses of our future work.
Acknowledgements
We would like to thank the anonymous reviewers for their inspiring questions and feedbacks for our work.
Impact Statement
This paper aims to advance the fields of Machine Learning and Robotics. While our study has numerous potential societal impacts, we believe none require specific emphasis in this context.
Appendix A Supplementary Materials
@@
A demo video, codes, and dataset for quantitative results and interpretation are available at https://sites.google.com/view/monet-lgc.
Appendix B Method Details
B.1 Perception Module in Details
@@
The perception module utilizes the Transformer encoder to generate a saliency map, which is then integrated with the global context of the input image $I$ through the self-attention mechanism. Following the description in (Dosovitskiy et al., 2020), the encoder network comprises: 1) a multi-head self-attention block (MSA), and 2) an MLP block, both equipped with layer normalizations (LN) and residual connections. After feature extraction by the CNN block (Eq. 13), the input embedding undergoes preprocessing (Eq. 14-15) before being fed into the Transformer encoder process (Eq. 16-18).
@@
Batch size 512
Total training iterations 650k
Optimizer Adam
Similarity factor $\kappa$ 0.5
Weight for the LGC loss term $\lambda_{LGC}$ 5e-4
Learning rate 3e-4
Learning rate scheduler LambdaLR
Scheduler factor 3e-4
Table 2: Hyperparameter configuration
Appendix C Experimental Details
C.1 Hyperparameter Setting
Table C.3 shows the hyperparameter setting for our experiments. The batch size is 512, and the total training iteration is 650k. We use Adam optimizer (Kingma & Ba, 2014) with an initial learning rate of 3e-4. For self-supervised learning, we set the similarity factor $\kappa$ to 0.5 to ensure that the LGC loss function conservatively identifies positive samples in the early phase of training.
C.2 Hardware System and Coarse Topology Map
Our platform consists of a 1/10 scale racing car chassis equipped with an embedded computer, Jetson Xavier NX, and a microcontroller, Arduino Nano. The Xavier NX receives front camera images with the Realsense D435i camera sensor and acquires range measurements using a 2D LiDAR sensor (Hokuyo UST-20LX). These measurements are utilized to estimate the current pose of the ego vehicle through onboard localization (Hess et al., 2016) in GPS-denied indoor environments. The Xavier NX then computes an ego-centric coarse topology map (Amini et al., 2019), which includes a highlighted routed map alongside an unrouted map, based on the ego vehicle’s pose and a globally routed path. This path is planned using the Dijkstra algorithm, utilizing a sparse topological roadmap of the indoor corridor environments. The Arduino Nano receives commands from either the Xavier NX or a human driver, converting them into PWM signals for the steering and speed control motors of the platform.
Figure 8: Hardware system setup.
C.3 Data Collection and Processing
While collecting data, we record camera images, topology maps, and corresponding command signals for steering and throttle control from the human driver. These control signals are normalized to a range of [-1, +1]. We collect data for a total of 2 hours, amounting to 88,326 pairs of sensory input and labels in the environments of Env. 1 and Env. 2. The data is split into training and validation sets at a ratio of $80:20$. The camera image is cropped to a size of $440\times 240$ pixels and then resized to $224\times 224$ pixels for use in our network. For data augmentation, we apply random image shifts and corresponding steering angle adjustments, as outlined in (Bojarski et al., 2016).
C.4 Perceptual and Behavioral Interpretability